Mary is twice as old as Jamie. Together their age add up to 54. How old is each of them?
Answer by Maths68(1474)

Let
Present age of Mary = m
Present age of Jamie = j

Mary is twice as old as Jamie.
m=2j........(1)
Together their age add up to 54.
m+j=54
Put the value of m in above equation
2j+j=54
3j=54
3j/3=54/3
j=18
Put the value of j in (1)
m=2(18)
m=36
Present age of Mary = m = 36 years old
Present age of Jamie = j = 18 years old